OpenStudy (lin.ivory):

Find the interest rate needed for an investment of 24000 dollars to grow to an amount of 34300 dollars in 8 years if the interest is compounded monthly.

OpenStudy (cwrw238):

sorry - i got called away formula is 34,300 = 24,000( 1 + r / 12)^(8*12) you have to solve this for r

OpenStudy (cwrw238):

can u continue?

OpenStudy (lin.ivory):

thats ok! thanks! I put 34300/24000 = (\[(1+ \frac{ r }{ 12 })^(96)\]

OpenStudy (lin.ivory):

to the power of 96

OpenStudy (lin.ivory):

but i dont know how to log both side? or get rid of the exponent 96

OpenStudy (cwrw238):

yes thats right

OpenStudy (cwrw238):

ok take logs of both sides ln (34300/24000) = 96 * ln (1 + r/12)

OpenStudy (cwrw238):

r will be in decimal form

OpenStudy (lin.ivory):

sorry but is it possible for you to explain ln? i dont quite understand the concept

OpenStudy (cwrw238):

wait we can do this another way not involving logs 34300/24000 = (1 + r/12)^ 96 (1 + r/12) = (34300/24000)^(1/96) ur calculator should be able to do this

OpenStudy (lin.ivory):

but the answer came out wrong?

OpenStudy (cwrw238):

i got 0.047 - or near that id have to do it again what did you get?

OpenStudy (cwrw238):

0.0447 or 4.47 %

OpenStudy (lin.ivory):

OOOooops!! i typed in the wrong decimal lol!

OpenStudy (cwrw238):

haha - we all do that sometimes

OpenStudy (lin.ivory):

thank you very much! :)

OpenStudy (cwrw238):

yw